### v1

# Stable Quantum Vortices in Lee-Huang-Yang Dipolar Superfluids

Abstract
The nucleation and dynamics of vortices in the quasi-two-dimensional rotating dipolar Bose-Einstein condensate are explored by taking into account the Lee-Huang-Yang (LHY) correction to the mean-field (MF) theory. Assuming approximate cancellation of the MF interactions, we focus on the formation of a pure LHY superfluid. The effect of rotational frequency $\Omega$ is investigated numerically by determining the corresponding number of stable vortices in the superfluid, together with the respective energy per particle $E$ and chemical potential $\mu$. The LHY superfluid provides a deep minimum of $E$ and $\mu$, indicating that it is a remarkably robust state of quantum matter. By fixing the LHY interaction strength, an exact single-vortex critical frequency is found, along with the respective chemical potential. A notable feature, observed when creating the LHY superfluid with fewer than five vortices, which is understood as being due to the superfluid’s nonlinearity and trapping aspect ratio, is the large frequency ranges admitting the production of two and four vortices, as compared to the small frequency ranges to obtain one and three vortices.
I Introduction
The experimental observation of dipolar Bose-Einstein condensates (BECs) in ultracold gases of 52Cr [3, 2, 1], 164Dy [5, 4] and 168Er [6] magnetic atoms has been a giant step forward in the studies of quantum matter [7, 8]. BECs of magnetic atoms exhibit long-range dipole-dipole interaction (DDI) between atoms, depending on the mutual orientation of atomic magnetic moments [9], in addition to nonlinear effects induced by the $s$-wave contact interactions between the atoms, including repulsion and attraction. On the other hand, techniques based on the Feshbach resonance can efficiently control the contact nonlinearity [10], enabling the control of the interplay between the nonlocal (DDI) and local (contact) interactions for performing various manipulations with the dipolar BECs. In particular, this setting was elaborated in Ref. [11] to control spatial separation of spin-coupled BECs. In this regard, it is relevant to note that tunable long-range interatomic correlations can be induced by laser-induced DDIs in BECs [12], creating a roton minimum in the excitation spectrum, reminiscent of the well-known properties of the strongly correlated superfluid represented by liquid helium II [13, 14]. Rotons are also well known as elementary excitations in dipolar BECs, see Refs. and references therein [15]-[18].
It is well known that the mean-field (MF) cubic self-attractions leads to the onset of the critical or supercritical collapse and eventual breakdown of the condensates, in the framework of the corresponding two- and three-dimensional (2D and 3D) Gross-Pitaevskii (GP) equations, respectively [19, 20, 21, 22]. Therefore, it is relevant to search for physically relevant mechanisms that can stabilize 2D and 3D self-trapped modes, such as bright solitons, solitary vortices, and others  [23, 24]. In this context, beyond-MF effects, which have been the subject of intensive experimental investigations [25, 26, 27], have shown a great potential. In this vein, it was predicted [28, 29] that the full stability of 2D and 3D self-trapped states in a binary BEC, in the form of quantum droplets (QDs), may be provided by the Lee-Huang-Yang (LHY) corrections to the GP equations [30, 31], which represent the effect of quantum fluctuations around the MF state. This result follows a previous suggestion [32], made in the context of the three-body Efimov physics [33], which pointed out that the BEC system with attractive two-body interactions could eventually be stabilized against the collapse by a repulsive three-body term [34], leading to the formation of self-bound boson droplets, termed boselets in Ref. [32], so that a trapping potential is not needed to hold the particles together.
In terms of the corresponding GP formalism, the correction to the MF approximation, as considered in Ref. [28], is represented by quartic self-repulsion term, which is usually much weaker than the cubic MF one. The LHY correction becomes significant in the special case when the intraspecies self-repulsion and interspecies cross-attraction terms in the corresponding system of nonlinearly coupled GP equations (mutuality fitted by means of the Feshbach resonance [35]) almost cancel each other for the symmetric state, with equal wave functions of the two components [28]. In particular, the “LHY superfluid” was predicted [36] and observed [37], in which the full cancellation of the MF fields makes the entire nonlinearity in the GP system represented solely by the LHY terms. In agreement with the predictions, the LHY corrections have made it possible to create robust QDs in binary BECs with the contact interactions [38, 39, 40, 41, 42], as well as in dipolar condensates dominated by the long-range magnetic DDIs (see theoretical predictions in Refs. [43, 44] and experimental realizations in Refs. [45, 46, 47]). In the latter case, the interplay between long-range interactions and the LHY effect also makes it possible to realize supersolids in the dipolar quantum gases [48, 49, 50], see also review [51]. The effects of quantum fluctuations may support supersolids in other settings as well [52, 53, 54, 55, 56, 57]. The role of the LHY corrections has also been demonstrated in experimental studies of the critical BEC temperature [58], quantum depletion  [59], excitation spectrum [60], and thermodynamic equation of state [61]. It is relevant to mention that, when investigating the universality of the contact interactions in the regime of observable beyond-MF effects, no measurable contributions from three-body interactions were found [62].
Quantized vortices, first predicted in the framework of GP equations for trapped Bose gases in Ref. [63], have been created in experiments using rotating magnetic traps or applying laser stirring to BEC above a certain critical rotational frequency. The observation of the vortices provides compelling evidence for the occurrence of superfluidity in BECs, sharing some features with liquid helium [64]. In dipolar Bose gases, quantized vortices have been intensively investigated, within the MF approximation, in the last decades up to now [65, 66, 70, 67, 68, 69, 71, 72, 73, 74, 75, 80, 81, 76, 77, 78, 79]. The structure and stability of vortices in dipolar BECs are strongly affected by the anisotropy of the DDIs and their strengths, with respect to the contact interactions [65, 70, 71, 82]. Static hydrodynamic solutions and their stability have been also explored for the rotating dipolar BECs in the Thomas-Fermi (TF) limit [83]. In this context, the structure of dipolar BECs was studied in the quasi-2D scenario with arbitrary orientations of the dipoles [84, 73]. The second-order-like phase transition of straight and helical vortex lines, occurring due to the effect of the dipoles’ orientation, has also been reported [68]. DDI strongly affects the number, structure, and stability of vortices. Further, the critical rotational frequency for the nucleation of vortices decreases with the increase of the DDI strength, while the contact interactions enhance the vortex stability [74] . The stabilization of 2D and 3D self-trapped (localized) QDs with embedded vorticity due to the action of the LHY effect in the binary BECs with contact interactions were predicted, respectively, in Refs. [85] and [86]. On the other hand, the QD solutions for axially symmetric vortex states in the single-component dipolar BEC were found to be fully unstable [87].
Although experiments have been carried out with 164Dy and 168Er dipolar condensates by taking into account the effects of quantum fluctuations, quantized vortices in LHY-dominated superfluids have not yet been explored in detail. The recent experimental observation of quantized vortices in a dipolar condensate of lanthanide atoms [88, 89] motivates further theoretical studies and new experiments with single- and multi-vortex states in one-component and binary rotating condensates [11] in the LHY-dominated regime. In light of the above, the main objective of the present paper is to report new results, based on systematic simulations, for the nucleation and dynamics of vortices in a one-component dipolar BEC in the LHY-dominated superfluid, featuring the interplay between dipolar and contact two-body interactions. The independent tunability of binary interactions and dipole-dipole interactions ensures that one may end up with a fluid completely dominated by quantum fluctuations alone.
The paper is organized as follows. In the next section, we formulate the single-component quasi-2D model including the DDI and LHY effect, basically following Ref. [11]. Systematic results, with the corresponding discussion, are reported in Section III. Section IV summarizes the paper with our main conclusions and perspectives.
II The GP system with the DDI and LHY terms
We consider a dipolar atomic BEC of $N$ atoms with mass $m$ and magnetic dipole moment $\mu_{M}$. The system is confined in a quasi-2D pancake-shaped harmonic-oscillator trap,
$V(\mathbf{r})\equiv\frac{m\omega_{\perp}^{2}}{2}\left(x^{2}+y^{2}+\lambda^{2}z^{2}\right),$} (1)
with a fixed large aspect ratio, $\lambda\equiv\omega_{z}/\omega_{\perp}=20$. In the rotating reference frame, the dynamics of the system is governed by the time-dependent GP equation for the wavefunction $\phi\equiv\phi({\mathbf{r}},t)$, which is subject to the unitary normalization [90, 91, 92]:
$\mathrm{i}\hbar\frac{\partial\phi}{\partial t}$ $=$ $\Big[-\frac{\hbar^{2}}{2m}\nabla^{2}+V(\mathbf{r})-\Omega L_{z}+g_{0}\left|\phi\right|^{2}+g_{{}_{\mathrm{LHY}}}\left|\phi\right|^{3}$ (2)
$+$ $\int U_{\mathrm{dd}}({\mathbf{r}}-{\mathbf{r}}^{\prime})\left|\phi^{\prime}\right|^{2}d^{3}{\mathbf{r}}^{\prime}\Big]\phi,$
where $L_{z}\equiv-\mathrm{i}\hbar\left(x\partial_{y}-y\partial_{x}\right)$ is the $z$-component of the angular momentum operator, with $\Omega L_{z}$ implying the rotation about the $z$-axis with angular frequency $\Omega$. The coefficient $g_{0}$ of the cubic MF term is given in terms of the two-body $s$-wave scattering length $a_{s}$, as $g_{0}={4\pi\hbar^{2}a_{s}N}/{m}$. The DDI is represented by the integral term in Eq.  (2), which refers to the quasi-2D geometry illustrated in Fig.  1 [9],
$U_{\mathrm{dd}}(\mathbf{r-r^{\prime}})\equiv\displaystyle\frac{\mu_{0}{\mu_{M}^{2}}N}{8\pi}\frac{\left(3\cos^{2}\varphi-1\right)}{|\mathbf{r-r^{\prime}}|^{3}}.$ (3)
The strength of the quartic nonlinear LHY term for the dipolar BEC is
$g_{\mathrm{LHY}}=\frac{64\hbar^{2}\sqrt{{\pi(a_{s}N)^{5}}}}{m}\left(\frac{2}{3}+\varepsilon_{\mathrm{dd}}^{2}\right),$ (4)
where $\varepsilon_{\mathrm{dd}}\equiv a_{\mathrm{dd}}/a_{s}$ is the ratio of the effective DDI scattering length,
$a_{\mathrm{dd}}\equiv m\mu_{0}\mu_{M}^{2}/(12\pi\hbar^{2}),$ (5)
and the $s$-wave two-body scattering length $a_{s}$ [7].
Figure 1: The representation of DDIs in the dipolar quasi-2D BEC in the coordinate space. Angle between vector $\mathbf{r}_{2}\mathbf{-r}_{1}$ and the $z$ axis is $\theta_{d}\approx 90^{\mathrm{o}}$. The variation of angle $\varphi$ between the magnetic dipoles and the $z$ axis alters the strength and sign of DDI, from repulsive to attractive.
II.1 DDI in the quasi-2D BEC
To define the DDI, we consider a pair of atoms placed at positions $\mathbf{r}$ and $\mathbf{r^{\prime}}$, with the respective vector $\mathbf{r-r^{\prime}}$ making angle $\theta_{d}$ with the $z$ axis, so that $\theta_{d}\approx 90^{\circ}$ for the oblate (pancake-shaped) configuration. As shown in Fig. 1, the atoms have their magnetic moments polarized, by a polarizing magnetic field, in the same direction, making angle $\varphi$ with the $z$ axis. Similar to Refs. [9, 11], in the present case, the tunability of the DDI is provided by time-dependent polarizing fields, which set the magnetic moments in rapid rotation around the $z$ axis. With the magnetic fields being a combination of the static (dc) component directed along the $z$ axis and the rapidly rotating (ac) one in the (x,y)(x,y)\ plane, one may neglect the motion of atoms driven by the ac field. With the time averaging performed over the rotation period, the corresponding 3D averaged DDI can be written in the scaled form as [9]
$\hskip-5.69046pt\left\langle V_{\mathrm{3D}}^{(d)}({\mathbf{r}-\mathbf{r^{\prime}}})\right\rangle={\mu_{0}\mu_{M}^{2}}\left(\frac{{3\cos^{2}\varphi}-1}{2}\right)\frac{1}{\left|{\mathbf{r}-\mathbf{r^{\prime}}}\right|^{3}},$ (6)
where $\mu_{0}$ is the free-space permeability, with the magnetic moment $\mu_{M}$ measured in units of the Bohr magneton $\mu_{B}$. As said above, for the oblate configuration adopted here, with the condensate distributed in the $(x,y)$ plane, we have $\theta_{d}\approx 90^{\circ}$, hence $\varphi$ remains an appropriate angular coordinate for tuning the DDI strength, making it repulsive in the interval of $0\leq\varphi<\varphi_{\mathrm{magic}}$, and attractive for $\varphi_{\mathrm{\ magic}}$ $\varphi\leq 90$} , where $\varphi_{\mathrm{magic}}\approx 54.7^{\circ}$ is the so-called “magic angle” [93]. DDI vanishes at $\varphi=\varphi_{\mathrm{magic}}$, irrespective of the magnitude of the magnetic moment.
Equations (2) and (4) demonstrate that the short-range binary interaction, long-range DDI, and LHY correction are interdependent, which can impede the search for stable configurations, particularly vortices. Therefore, it may be helpful to develop a model exhibiting three independent interactions, which do not impact on each other. To this end, it is desirable to have the $g_{\mathrm{LHY}}$ coefficient which is affected not by the DDI orientation, but solely by the corresponding DDI scattering length (5). Therefore, a simple strategy to increase the impact of the LHY contribution is to assume a scenario in which one can adjust the contact and dipolar interactions, so that they effectively cancel out each other. To this aim, starting with a repulsive contact interaction, the long-range DDI should be tuned by adjusting the orientation angle $\varphi$ in Fig. 1, making it attractive. Provided that the strength of the repulsive contact interaction is adjusted to stay within the possible range of the attractive DDI strength, it may be possible to identify the setting in which the LHY effect dominates the condensate. Similarly, with repulsive DDI, one can manipulate the contact interaction, tuning it to the attraction sign, again designing the LHY-controlled setting, with the dipole-dipole and contact MF interactions balancing each other.
Appropriate dimensionless parameters can be defined in terms of a reference frequency, which we set as $\omega_{\perp}=\omega_{x}=\omega_{y}$, see Eq. (1), and the corresponding oscillator length, $\ell_{\perp}=\sqrt{\hbar/(m\omega_{\perp})}$. To cast Eq. (2) in the normalized form, we first rescale the time and coordinate variables as $\bar{t}\equiv\omega_{\perp}t$, $\bar{\mathbf{r}}\equiv\mathbf{r}$. Accordingly, we define $\bar{a_{s}}=a_{s}/\ell_{\perp}$, $\bar{a}_{\mathrm{dd}}=a_{\mathrm{dd}}/\ell_{\perp}$, with the wavefunction and system’s energy rescaled as $\bar{\phi}=\ell_{\perp}^{3/2}\phi$ and $\bar{E}=\hbar\omega_{\perp}E$. Next, we drop the overhead bars, assuming that all the variables and parameters are dimensionless. This transformation casts the GP equation (2) into the normalized form, with $\hbar=m=1$ and redefined coefficients.
In our approach, we assume a strong harmonic trap confinement in the $z$-direction, allowing the 3D ground-state wavefunction be expressed by the ansatz $\phi(\mathbf{r},t)\equiv{(\lambda/\pi)^{1/4}}\exp\left({-\lambda z^{2}}/{2}\right)\psi(\boldsymbol{\rho},t),$ where $\psi\equiv\psi(\boldsymbol{\rho},t)$ is the effective 2D MF wavefunction in polar coordinates, where $\boldsymbol{\rho}\equiv(x,y)=(\rho\cos\theta,\rho\sin\theta)$. The factorized ansatz is substituted in Eq. (2), which is then multiplied by $\phi(z)\equiv$ ${(\lambda/\pi)^{1/4}}\exp\left({-\lambda z^{2}}/{2}\right)$ and integrated over $z$ [94]. The corresponding procedure for the 2D reduction of the DDI term is elaborated in detail in Refs. [75, 11], in which the effective 2D DDI was derived by applying the convolution theorem and using the 2D Fourier transform for the product of the DDI kernel and density, as $\int d{{\ \boldsymbol{\rho}}^{\prime}}V^{(d)}({{\boldsymbol{\rho}}-{\boldsymbol{\ \rho}}^{\prime}})|\psi({\boldsymbol{\rho^{\prime}}})|^{2}=\mathcal{F}_{\mathrm{2D}}^{-1}[\widetilde{V}^{(d)}(k_{x},k_{y})\widetilde{n}(k_{x},k_{y})]$, where $\mathcal{F}_{\mathrm{2D}}^{-1}$ is the operator of the inverse Fourier transform. As a result, the DDI kernel can be expressed as the combination of two terms in the 2D momentum space, $\mathbf{k_{\rho}}=(k_{x},k_{y})=(k_{\rho}\cos\theta_{k},k_{\rho}\sin\theta_{k})$, for a fixed orientation $\varphi$ of the magnetic dipoles (Fig. 1), accounting for the interactions perpendicular and parallel to the direction of the dipole moments:
$\widetilde{V}_{\perp}^{(d)}(k_{x},k_{y})$ $=$ $2-3\sqrt{\frac{\pi}{2\lambda}}k_{\rho}\exp\left(\frac{k_{\rho}^{2}}{2\lambda}\right)\mathrm{erfc}\left(\frac{k_{\rho}}{\sqrt{2\lambda}}\right),$
$\widetilde{V}_{\parallel}^{(d)}(k_{x},k_{y})$ $=$ $-1+3\frac{k_{x}^{2}}{k_{\rho}}\sqrt{\frac{\pi}{2\lambda}}\exp\left(\frac{k_{\rho}^{2}}{2\lambda}\right)\mathrm{erfc}\left(\frac{k_{\rho}}{\sqrt{2\lambda}}\right),$ (7)
where erfc($\xi$) is the standard complementary error function. With all directions being possible in the $\left(x,y\right)$ plane, the angular averaging amounts to $k_{x}^{2}=k_{\rho}^{2}\cos^{2}\theta_{k}\rightarrow k_{\rho}^{2}/2$, so that the effective total 2D DDI kernel can be written in the momentum space as a function of $k_{\rho}$ and $\varphi$, viz, $V_{\varphi}(k_{\rho})\equiv\widetilde{V}^{(d)}(k_{x},k_{y})$:
$V_{\varphi}(k_{\rho})$ $=$ $\cos^{2}\!\varphi\widetilde{V}_{\perp}^{(d)}(k_{x},k_{y})+\sin^{2}\!\varphi\widetilde{V}_{\parallel}^{(d)}(k_{x},k_{y})$
$=$ $\frac{3\cos^{2}\!\varphi-1}{2}\left[2-3\sqrt{\frac{\pi}{2\lambda}}\;k_{\rho}\mathrm{e}^{\frac{k_{\rho}^{2}}{2\lambda}}\mathrm{erfc}\left(\frac{k_{\rho}}{\sqrt{2\lambda}}\right)\right].$
Thus, we have the effective GP equation given by
$\mathrm{i}\frac{\partial\psi}{\partial t}$ $=$ $\biggr[-\frac{\nabla_{\rho}^{2}}{2}+\frac{\rho^{2}}{2}-\Omega L_{z}+g|\psi|^{2}+\eta\left|\psi\right|^{3}\biggr]\psi$
$+$ $g_{\mathrm{dd}}(\varphi)\int\frac{d^{2}\mathbf{k}_{\rho}}{(2\pi)^{2}}e^{-i\mathbf{k}_{\rho}.\boldsymbol{\rho}}\widetilde{n}(\mathbf{k}_{\rho},t)h_{\mathrm{2D}}\biggr(\frac{k_{\rho}}{\sqrt{2\lambda}}\biggr)\psi,$
where $g_{\mathrm{dd}}(\varphi)\equiv(g_{\mathrm{dd}}/2){(3\cos^{2}\!\varphi-1)}$. Here, the coefficient $\eta$, which is defined below by Eq. (15), corresponds to the reduced LHY dimensionless coefficient $g_{\mathrm{LHY}}$ given by Eq. (4) for the dipolar system. The other factors in Eq. (II.1) are [95]
$\widetilde{n}(\mathbf{k}_{\rho},t)$ $=$ $\int d^{2}\boldsymbol{\rho}e^{i\mathbf{k}_{\rho}.\boldsymbol{\rho}}|\psi|^{2},$
$h_{\mathrm{2D}}(\xi)\Big|_{\xi\equiv\frac{k_{\rho}}{\sqrt{2\lambda}}}$ $=$ $\left[2-3\sqrt{\pi}\xi\exp(\xi^{2})\text{erfc}(\xi)\right].$ (10)
[…]
The total energy corresponding to Eq. (II.1) can also be written as [97]
$E$ $=$ $\int\!\!dxdy\left\{\frac{1}{2}\left[\left|\mathrm{i}\partial_{x}\psi-\Omega y\psi\right|^{2}+\left|\mathrm{i}\partial_{y}+\Omega x\psi\right|^{2}\right]\right$.
$+$ $\left.\frac{(1-{\Omega}^{2})(x^{2}+y^{2})}{2}|\psi|^{2}+\frac{g}{2}|\psi|^{4}+\frac{2}{5}\eta|\psi|^{5}\right\}$
$+$ $\frac{g_{\mathrm{dd}}(\varphi)}{2}\int\frac{d^{2}\mathbf{k}_{\rho}}{(2\pi)^{2}}e^{-i\mathbf{k}_{\rho}.\boldsymbol{\rho}}\widetilde{n}(\mathbf{k}_{\rho},t)h_{\mathrm{2D}}\biggr(\frac{k_{\rho}}{\sqrt{2\lambda}}\biggr)|\psi|^{2}\Bigg\}.$
In the limit of extremely tight confinement in the transverse direction, the effective nonlinearity in the LHY-amended GPE is different from the quartic term [29]. Nevertheless, one may use the 2D equations with the quartic term, such as Eq. (II.1), in relevant experimental settings with moderately tight confinement [98].
II.2 The LHY correction for the quasi-2D BEC
The LHY corrections can be taken into account in the usual MF 3D formalism similar to how it was done in Ref. [36], where the correction to the ground-state energy density of a homogeneous weakly repulsive Bose gas with contact interactions, is
$\mathcal{E}_{\mathrm{LHY}}$ $=$ $\frac{\hbar^{2}}{15m}{256\sqrt{\pi(|a_{s}|n)^{5}}}$
$=$ $\frac{\hbar\omega_{\perp}}{\ell_{\perp}^{3}}\frac{256\sqrt{\pi}}{15}{{\left(N\frac{|a_{s}|}{\ell_{\perp}}|\psi(x,y)|^{2}\sqrt{\frac{\lambda}{\pi}}e^{-\lambda z^{2}}\right)^{5/2}}}.$
The corresponding LHY corrections are given by
$H_{\mathrm{LHY}}\Psi=\left(\frac{128\sqrt{\pi}}{3}\sqrt{\frac{N|a_{s}|^{5}}{\ell_{\perp}^{5}}}\right)|\psi(x,y)|^{3}|\phi(z)|^{3}\Psi.$ (13)
Multiplying this term by the complex conjugate of $\phi(z)$, and integrating out the $z$-component, we obtain the following expression for the scaled LHY correction term,
$H_{\mathrm{LHY}}\psi=\frac{128}{3}\sqrt{\frac{2N^{3}}{5}}\left(\frac{\lambda^{3}}{\pi}\right)^{1/4}\left({\frac{|a_{s}|}{\ell_{\perp}}}\right)^{5/2}|\psi|^{3}\psi,$ (14)
which provides the scaled coefficient corresponding to coefficient $g_{\mathrm{LHY}}$ in Eq. (4), in the absence of DDI. In the case of GP equation for the dipolar gas, the coefficient $g_{\mathrm{LHY}}$ needs to be multiplied by an additional factor, which depends on the DDI strength, as shown in Refs. [90, 91, 92, 99], thus giving rise to the interdependence between the LHY correction, contact interactions, and DDI. Including the DDI, we thus derive the effective coefficient for the scaled quasi-2D LHY correction as
$\eta=\frac{64\sqrt{2}}{\sqrt{5\sqrt{\pi}}}\left({\sqrt{\lambda}N}\right)^{\frac{3}{2}}\left(\frac{|a_{s}|}{\ell_{\perp}}\right)^{\frac{5}{2}}\left(\frac{2}{3}+\varepsilon_{\mathrm{dd}}^{2}\right).$ (15)
It appears in Eq. (4), for a quasi-2D dipolar condensate with aspect-ratio parameter $\lambda$, that was defined above in Eq. (1).
II.3 The pure LHY nonlinearity in the quasi-2D GP equation
In the above derivation of Eq.(15) it is adopted that the dipoles are aligned in the same direction (Fig. 1), implying that the DDI is maximized at $\varphi=0^{\circ}$ [90]. As also seen in Fig. 1, the effective DDI strength can be tuned by varying $\varphi$ from 0 (maximum positive) to $\pi/2$ (maximum negative), being zero for $\varphi=\varphi_{\mathrm{magic}}$. To take advantage of the tunability of DDI  [9], we keep $\varphi$ as an adjustable parameter. For this purpose, the LHY parameter (15) is redefined, as $\eta\equiv\eta(\varphi)$, with
$\varepsilon_{\rm{dd}}\equiv\varepsilon_{\mathrm{dd}}(\varphi)\equiv\frac{a_{\mathrm{dd}}}{a_{s}}\left(\frac{3\cos^{2}\varphi-1}{2}\right).$ (16)
Thus, the effective LHY coefficient is affected by two adjustable parameters: $a_{s}$ and the angle $\varphi$. Independently of the signs of the contact and dipolar interactions, the LHY contribution is positive. The possibility to enhance the LHY contribution as compared with the main nonlinear terms, is to make use of both control parameters, so that one of the interactions is attractive, with the other being repulsive. The ideal condition for a pure LHY superfluid, is the overall total cancellation of the MF and dipolar terms. In the framework of a more realistic experimental setup, one can consider the situation in which we have an effective reduction of the two main nonlinear terms such that the fluid’s nonlinearity is dominated by the quantum fluctuations. In the following section, we consider the ideal condition in which we can effectively remove the MF terms, viz., DDI and contact ones, in Eqs. (II.1) and (II.1). The result is the following stationary 2D GP equation for the quantum superfluid (QS) with chemical potential $\mu_{\mathrm{QSF}}$, and the respective energy per particle:
$\hskip-14.22636pt\mu_{QS}\psi$ $=$ $\biggr[-\frac{1}{2}\left({\partial_{x}^{2}}+{\partial_{y}^{2}}\right)+\frac{x^{2}+y^{2}}{2}-\Omega L_{z}+\eta(\varphi)\left|\psi\right|^{3}\biggr]\psi,$
${E_{\mathrm{QS}}}$ $=$ $\int\!\!dxdy\Bigg\{\frac{1}{2}\left[\left|\mathrm{i}{\partial_{x}}-\Omega y\psi\right|^{2}+\left|\mathrm{i}{\partial_{y}}+\Omega x\psi\right|^{2}\right]$
$+$ $\frac{(1-{\Omega}^{2})(x^{2}+y^{2})}{2}|\psi|^{2}+\frac{2}{5}\eta(\varphi)|\psi|^{5}\Bigg\}.$
This possibility to effectively cancel the local and nonlocal MF nonlinear terms are somewhat similarity to that proposed in Ref. [36], where a binary Bose mixture without DDI and rotation was considered, therefore our studies of on rotational properties of LHY superfluid, following below in section III.3, may be accordingly applied to the setting addressed in Ref. [36]. In addition, it is relevant to mention that, for the 2D BEC trapped in the harmonic-oscillator potential, the critical rotational frequency $\Omega_{c}$ for a single vortex production has been analytically determined in Ref. [100], assuming the strong cubic nonlinearity obeying the Thomas-Fermi limit. In the present notation, the corresponding expression is
$\Omega_{c}\approx\frac{5}{2}\ln\left(0.67\sqrt{2\mu}\right).$ (18)
In the case of more general nonlinear interactions, such as the present case, which includes the DDI and LHY terms, the result is expected to be similar, as the essential differences originating from the nonlinear interactions, may be absorbed by the change of the chemical potential.
III Numerical results and discussion
The main objective of this work is to explore the nucleation and dynamics of vortices in the BEC dominated by the LHY nonlinearity, assuming the effective mutual cancellation of the MF contact interactions and DDIs. In subsection A, the results are produced in the absence of the LHY term, aiming to verify the rotation effect in the usual situation, where the MF nonlinearity is taken into account. These results, which are consistent with previous investigations, serve as a benchmark for the comparison with the cases where the LHY term is the dominant one. In that subsection, the explicit dependence on the orientation angle $\varphi$ is dropped in the DDI and LHY parameters, assuming that $\eta\equiv\eta(\varphi)$ is given by Eq. (16) and ${g}_{\mathrm{dd}}\equiv g_{\mathrm{dd}}(\varphi)$. Unless explicitly stated otherwise, this implies $\varphi=0^{\circ}$. In particular, the pure LHY superfluid considered in subsection C relies on the assumption of the effective cancellation of the MF contact and DDI terms, making it relevant to keep the $\varphi$ dependence of the $\eta$ and $g_{\mathrm{dd}}$ coefficients.
III.1 Vortex modes in the absence of the LHY term
The numerical procedure to solve the 2D GP equation starts with the split-step Crank-Nicholson method. Its advantage is that the non-derivative nonlinear and linear terms can be treated very accurately, to secure the precision and robustness of the numerical scheme. In the present study, we have used the numerical code with stepsizes $\Delta x=$ $\Delta y=0.04$ and $\Delta t=0.001$ [94, 96, 101].
Figure 2: (Color online) Density profiles $|\psi|^{2}$ projected onto the 2D plane, displaying vortex lattices with the nonlinearity represented solely by the MF contact interactions with $g=100$. The rotation frequences $\Omega$ (in units of $\omega_{\perp}$) are indicated in the panels, with the density levels coded by the by color bars. The $x$ and $y$ coordinates (in units of $\ell_{\perp}$) cover the interval [10,+10][-10,+10].
The numerical results are exemplified by the vortex formation in two cases when the MF nonlinear terms in the GP equation represent the contact and dipolar interactions. The obtained results are consistent with previously reported ones, such as those for the rotating systems with the pure contact nonlinearity [63], and for the 3D dipolar rotating system with a high value of the aspect ratio [102]. Here, it is not only illustrative consider these results, which are also useful to estimate effects produced by the MF terms in comparison to those produced by the LHY nonlinearity. For this purpose, we keep the same DDI and contact-interaction parameters throughout the present study. Figure 2 shows how the vortices emerge following the variation of the rotation frequency $\Omega$ in the usual nonlinear MF system, with $g=100$, in the absence of the DDI and LHY terms ($g_{\mathrm{dd}}=\eta=0$) in Eq. (II.1). Consistent with the critical frequency (18) for the single vortex in the 2D BEC trapped in the harmonic-oscillator potential with frequency $\omega_{\perp}$ (where $\mu\approx 2.58$ is the scaled chemical potential), it is observed that the generation of vortices starts slightly below $\Omega\approx 0.6$, with the number of vortices $N_{v}$ increasing almost linearly in the interval of $0.7<\Omega<0.9$. In Fig 3, our focus is on the generation of vortices with the nonlinearity represented solely by the DDI term, so that $(g,g_{\mathrm{dd}},\eta)=(0,100,0)$. Comparing the results displayed in Figs. 2 and  3, one can understand the efficiency of the respective parameters for the control of the production of vortices. These two cases, presented in Figs. 2 and 3, in the absence of the LHY nonlinearity, will be useful for comparison with the impact of the LHY term on the vortex production, which is the subject of the next subsection.
Figure 3: (Color online) Density profiles $|\psi|^{2}$ projected onto the 2D plane, displaying vortex lattices with the nonlinearity represented solely by DDI with $g_{\mathrm{dd}}=100$. The rotation frequencies $\Omega$ are indicated in the panels, with the density levels coded by the color bars. The spatial domain and units are the same as in Fig. 2.
III.2 Vorticity in dipolar BEC with contact and LHY nonlinearities
The effect of LHY correction to the MF GP equation is estimated in this section by assuming $\eta=200$ in Eq.(II.1), with $g=100$ and without dipolar interactions. By strictly considering the expression (15), we should be aware that any specific value assumed for $\eta$ will depend on $a_{s}$ and $a_{\mathrm{dd}}$. In particular, when $a_{s}=0$, the quantum fluctuations become zero, such that $\eta\neq 0$ with $a_{s}=0$ has no interest in our present analyses. However, we can verify rotational effects due to quantum fluctuation in a non-dipolar BEC system. For such a case, by assuming $g_{\mathrm{dd}}=0$ in Eq.  (II.1), we first consider the results obtained in Fig. 2 with $g=100$ to estimate how strong the parameter $\eta$ of the quantum fluctuations could be assumed to reach some visible effect in the results. By looking for that, we have fixed $\eta=200$ in our following studies, with the results being presented in Fig. 4. The quantum fluctuation contribution to the results, as expected, is quite small when compared with the MF nonlinear one, such that it starts to be noticeable only with $\eta=200$ at $\Omega$ between 0.5 and 0.6. This can be verified in our results of Figs. 2 and  4. The effect of quantum fluctuations with such $\eta$ can also be noticed for large rotations, close to $\Omega=0.99$.
Figure 4: (Color online) Density profiles $|\psi|^{2}$ projected onto the 2D plane, displaying vortex lattices with nonlinearity provided by contact and LHY interactions, with $g=100$ and $\eta=200$. The frequencies are indicated inside the panels, with the respective density levels given by color bars. The spatial domain and units are the same as in Fig. 2.
With the given results in Fig. 5, we are further verifying the significance of the LHY correction with $\eta=200$. In this case, by considering both contact and dipolar nonlinearities together with the quantum fluctuations, one should notice that LHY quantum fluctuation with $\eta=200$ remains producing a small effect. The main change in the observed number of vortices in Fig. 5 is due to the cumulative effects related to contact and dipolar interactions, as compared with previous results shown in Figs. 2 and 3. Therefore, a possible way to enhance the effect of LHY correction, in comparison with other nonlinear effects, relies on playing with the signals of dipole-dipole and contact interactions, which can be performed by noticing that the strength of the LHY term, in the case of dipolar BECs, is only affected by the square of $\varepsilon_{\mathrm{dd}}$, as shown by Eq. (16). Therefore, one can maximize or minimize the DDI effect just by controlling the $\varphi$ dependence of $\eta$. This is being considered in the next section III.3 as a possible mechanism to generate a pure-LHY dipolar superfluid.
Figure 5: (Color online) Density profiles $|\psi|^{2}$ projected onto the 2D plane, displaying vortex lattices with nonlinearity provided by contact, dipolar and LHY interactions, with $g=100$, $g_{dd}=100$, and $\eta=200$. The frequencies are indicated inside the panels, with the respective density levels given by color bars. The spatial domain and units are the same as in Fig. 2.
III.3 The generation of vortices in the pure-LHY superfluid
The scenario where the LHY correction is the only source of the nonlinearity in the single-component dipolar BEC is explored in this subsection. It should be pointed out at this juncture that nonlinearity solely due to LHY was previously considered in the two-component BEC mixture with contact interactions [36, 37]. When the intra-species repulsion and inter-species attraction are precisely balanced, the MF contributions can be minimized or exactly canceled. To provide the approximate cancellation of the nonlinear MF local and nonlocal terms in the framework of the GP equation, we here consider the quasi-2D single-component dipolar BEC system introduced in the previous section. In this scenario, one can design the DDI intensity, by suitably adjusting the angle $\varphi$ of the dipoles with respect to axis $z$ (as represented in Fig. 1), and the contact interaction, through the scattering length $a_{s}$, by using the Feshbach resonance mechanism. Following that, within some approximation, we are left solely with the LHY nonlinearity, as postulated by Eq. (II.3), where the quintic term represents the LHY nonlinearity, with no explitic contribution from the contact and dipolar interactions.
Figure 6: (Color online) Density profiles $|\psi|^{2}$ projected onto the 2D plane, which display vortex lattices supported solely by the LHY nonlinearity, with $\eta=200$ in Eq. (II.1). The rotation frequencies $\Omega$ are indicated in the panels, with the respective density levels defined by the color bars. The spatial domain and units are the same as in Fig. 2.
The rotating dipolar LHY superfluid is exemplified by setting $\eta=200$ in Eq. (II.3). For this dipolar BEC system, the values of $a_{s}$ and $a_{\mathrm{dd}}$, which are included in expression (15), are not critical, provided that they are tuned to secure the approximate cancellation of the local and nonlocal MF terms. By changing the rotation frequency $\Omega$ from 0, by increments of $\Delta\Omega=0.1$, we have observed no vortex produced for $\Omega\leq 0.6$ (in units of $\omega_{\perp}$), with a pair of vortices being verified at $\Omega=0.7$. Next, two and four pairs ($N_{v}=4$ and $8$) are observed at $\Omega=0.8$ and $\Omega=0.9$, respectively. These results are shown in Fig. 6. In more detail, we explored the values of $\Omega$ to produce the first four vortices, with results for the density plots and corresponding phases presented in Fig. 7. In this regard, it was verified that, to produce one or three stable vortices, careful fine tuning of $\Omega$ is required, with the values of $\Omega$ being, respectively, $0.6401$ and $0.768$. These findings reveal criticality in the states concerning $\Omega$, for relatively small number of produced vortices, $N_{v}$. More explicitly, at $\Omega=0.6400$, no vortex can be found, whereas two vortices are generated at $0.6410\leq\Omega\leq 0.767$, as shown in Fig. 7(b). Next, by slightly increasing the rotation frequency to $\Omega=0.768$, a state with three-vortex is obtained, which persists in a very narrow interval, $0.768\leq\Omega 0.778$. A small variation in $\Omega$ can transform the three-vortex state into a two- or four-vortex one (at $\Omega\leq 0.767$ and $\Omega\geq 0.779$, respectively. The four-vortex state is produced in a broader interval, $0.779\leq\Omega\leq 0.881$, with Fig. 8(d) displaying the result for $\Omega=0.779$. This behavior, with vortex configurations being more stable with $2$ and $4$ vortices than with $1$ and $3$ vortices, is apparently a characteristics of the pure LHY superfluid in our quasi-2D approach, which is represented by the quartic nonlinearity in the GP equation with a strong pancake-like aspect ratio.
Figure 7: (Color online) For the LHY superfluid with $\eta=200$ at four values of $\Omega$ (in units $\omega_{\perp}$), the density profiles $|\psi|^{2}$, with respective phases patterns shown in two upper rows, with the corresponding 3D representations drawn in the lower rows part of the figure. Panel (a) pertains to the critical frequency, $\Omega_{c}=0.6401$, for the creation of the single-vortex configuration. Coordinates $(x,y)$ cover intervals [5,+5][-5,+5] (in units of $\ell_{\perp}$).
Figure 8: (Color online) In the case when the nonlinearity is represented solely by the LHY term, we show the long-time evolution of the angular momentum $\langle L_{z}\rangle$ (in units of $\hbar$), corresponding to the production of one to four vortices, as indicated in panel (a). In panel (b) we display the same case as in (a) for a much shorter time interval, $t<70$ (in units $\omega_{\perp}^{-1}$). The respective values for the rotation velocity $\Omega$ are given in the caption of Fig. 7. The four established values, $\langle L_{z}\rangle=1,1.75,2.35$, and $2.85$, agree with the ones reported in Ref. [63] for the case of pure contact interactions.
Our results for the pure-LHY nonlinearity are consistent with results reported in Ref. [63], in which only the contact interactions have been considered, and the predictions were produced for rotating states within a more general 3D framework. For the first four stable vortices shown in Figs. 7, with $\Omega=0.6401,0.641,0.768$, and $0.779$, the respective values of the system-averaged angular momentum per particle is $\langle L_{z}\rangle=1,1.75,2.35$ and $2.85$. These results are analyzed in Fig. 8, considering the respective time evolution of $\langle L_{z}\rangle$. Figure 8 shows how far one needs to evolve $\langle L_{z}\rangle$ to obtain stable vortex states. In addition, as the single-vortex configuration shown in (a1) of Fig. 7 can only be obtained after fine-tuning the rotational frequency within four decimal digits, we can identify the critical rotation frequency for the production of a single vortex in the pure-LHY case as $\Omega_{c}=0.6401$. The corresponding numerical value of the chemical potential is $\mu_{\mathrm{QS}}=1.78$. This result is close enough to the one that can be extracted from the TF relation (18), $\mu\approx 1.86$, for the same $\Omega_{c}$.
Usually, the vortex patterns are made energetically preferable states with the help of the trap symmetry and boundary conditions. As shown in Refs. [104, 103, 63], for contact interactions, the critical rotation frequencies $\Omega_{cn}$, which are necessary to attain stable states with $n$ vortices are determined by the MF two-body interaction. Therefore, for vortices supported by the pure LHY nonlinearity, such critical frequencies and the corresponding stability may be estimated by means of the respective coefficient $\eta$, given in Eq. (15).
Figure 9 summarizes the results for $N_{v}$ versus $\Omega$, varying $\Omega$ up to $0.99$, for the solely-LHY nonlinearity, with strength $\eta=200$. In this case, the largest number of vortices found for $\Omega=0.99$ is $N_{v}=20$. The figure displays the range intervals of frequencies $\Delta\Omega$ for the given specific values of $N_{v}$. Notable are the cases with $N_{v}=2,4$ and $8$, which have larger $\Delta\Omega$ than the other cases. Also, noticiable is the case with $N_{v}=1$, which has $\Omega_{c}=0.6401$ with $\Delta\Omega<0.001$ (In this regard, see also the upper row of Fig.7). Inside Fig. 9, we have also included density plots obtained for $N_{v}=5,6,7$ and $10$, to complement the ones previously shown for $N_{v}\leq 4$.
Figure 9: The dependence of the number of vortices, $N_{v}$, on the rotation frequency $\Omega$ (in units of $\omega_{\perp}$) for the pure LHY case, with $\eta=200$ (see Eq. (15)). The position of the single-vortex critical frequency, $\Omega_{c}=0.6401$, is indicated. For even numbers $N_{v}\geq 2$, the existence ranges of $\Omega$ are indicated too. The density profiles for $N_{v}=5,6,7,$ and $10$ are plotted in the panel.
III.4 The effect of the rotational frequency
For the usual BEC system trapped in the harmonic oscillator potential, with the cubic MF nonlinearity, the rotating states were first predicted in Ref. [63] in the framework of a more general 3D framework, considering the generation of vortices, with the corresponding angular momentum being a function of the rotation velocity $\Omega$. The previous section, confirms the occurrence of the same kind of vortex-generating mechanism in the LHY-dominated condensate, as in the case of the one with the MF nonlinearity. By adding the rotation to the system, all vortices are produced with the same vorticity sign, without vortex-antivortex annihilation.
Table 1: The effect of rotation frequency $\Omega$ on chemical potential ($\mu$), total energy per particle ($E$), number of visible vortices ($N_{v}$), and angular momentum per particle, $\left\langle L_{z}\right\rangle$ for five different sets of parameters considered in this work. Case I (the pure MF contact interaction, with $g=100$); case II (the pure DDI, with $g_{\mathrm{dd}}=100$); case III (with the contact-MF and LHY terms, without DDI, for $g=100$ and $\eta=200$); case IV (the nonzero contact-MF, DDI, and LHY terms, with $g=100$, $g_{\mathrm{dd}}=100$, and $\eta=200$); and case V (the pure LHY nonlinearity, with $\eta=200)$. In the pure-LHY superfluid, one should observe that the critical frequencies $\Omega_{c}$ for producing one and two vortex states are about the same within two decimal digits (being, respectively, $0.6401$ and $0.6410$). It implies that a single-vortex production is numerically quite challenging. In this table, the corresponding values for $N_{v}=1$ are: $\Omega=0.6401$, $\mu=1.78$, $E=1.21$, and $\langle L_{z}\rangle=1$.
$\Omega$ $\mu$ ($\hbar\omega_{\perp}$) $E$ ($\hbar\omega_{\perp}$) $N_{v}$ $\langle L_{z}\rangle$ ($\hbar$)
I II III IV V I II III IV V I II III IV V I II III IV V
0.0 2.96 4.61 3.40 5.62 2.06 2.05 3.10 2.30 3.76 1.40 0 0 0 0 0 0 0 0 0 0
0.4 2.96 4.61 3.40 5.62 2.06 2.05 3.10 2.30 3.76 1.40 0 0 0 0 0 0 0 0 0 0
0.5 2.96 4.12 3.40 4.99 2.06 2.05 2.68 2.30 3.18 1.40 0 4 0 4 0 0 2.76 0 2.99 0
0.6 2.59 3.80 2.83 4.37 2.06 1.76 2.42 1.84 2.77 1.40 2 4 4 8 0 1.69 3.06 2.72 4.68 0
0.7 2.24 3.07 2.50 3.54 1.61 1.48 1.89 1.58 2.16 1.09 4 8 4 10 2 2.85 4.91 3.05 6.13 1.71
0.8 1.62 2.22 1.69 2.51 1.23 0.99 1.23 0.96 1.26 0.80 8 12 8 14 4 4.34 6.98 5.00 8.44 2.88
0.9 0.38 0.24 0.47 0.04 0.55 -0.05 -0.47 -0.07 -0.03 0.24 14 22 14 28 8 7.19 11.38 8.10 14.12 4.53
0.99 -3.81 -6.06 -4.30 -7.37 -1.79 -3.99 -6.37 -4.51 -7.76 -1.89 42 68 48 80 20 20.75 30.94 22.73 36.53 11.99
The findings for the vortex configurations are summarized in Table 1 for all five nonlinearity cases that we have considered. In the table, we report the numerical results obtained for the chemical potential, total energy per particle $E$, and number of visible vortices $N_{v}$, for specific values of the rotation frequency $\Omega$, which are given with one decimal digit. For the five cases being considered, the corresponding behaviors of $E$ and $\mu$, as functions of $\Omega$, are presented in two panels of Fig. 10. The energy cost of generating vortices can be verified from the results provided in Table 1. In particular, by focusing on case V of the table, where we have only the LHY nonlinear term, the energy cost of producing a pair of vortices, as compared with the no-rotating case ($E(\Omega=0)$), is $\approx 22\%$ at $\Omega=0.7$. On the other hand, when we have only the MF contact nonlinearity (column I in the table), the cost is $\approx 14\%$ at $\Omega=0.6$. This indicates that the vortices are more likely to be produced in the presence of the pure MF contact nonlinearity than in the case of the pure LHY nonlinearity.
Figure 10: (Color online) Energy $E$ per particle (a) and the corresponding chemical potential $\mu$ (b) are shown as functions of $\Omega$, for five different combinations of the interactions considered in Table 1 (as indicated in the panels). The energy and frequency units are, respectively, $\hbar\omega_{\perp}$ and $\omega_{\perp}$.
Figure 11: (Color online) The linear increase of the angular momentum per particle $\langle L_{z}\rangle$ (in units of $\hbar$), associated with the change in the number of vortices, $N_{v}$, is confirmed for all the cases presented in Table 1. As shown, the results are close to the linear behavior (solid line) predicted by Feynman [13]. The parameters are indicated in the panel, where QS stands for the solutions corresponding to Eq. (II.3), pertaining to the pure-LHY nonlinearity.
Similarly, as shown in Ref. [100] for the contact MF nonlinear interaction, the critical frequency necessary to produce the first vortex is given by the approximate analytical expression (18). Essentially, such an approximate expression in terms of the chemical potential is expected to be similar for all five cases considered in Table 1. Given the exact numerical results for $\mu(\Omega)$, when starting to deviate from the no-rotation case, $\mu(\Omega=0)$, we can obtain the corresponding single-vortex critical frequencies. Roughly, these critical frequencies can be numerically extracted from the results provided in Fig. 10, as well as from Table 1.
We conclude this section by showing that all the obtained results can be well fitted by the Feynman prediction of the linear behavior of the angular momentum per particle versus the number of vortices [13], with the results provided in Fig. 11.
IV Conclusion
In the present work, considering the rotating dipolar BEC, we investigate the impact of quantum fluctuations, represented by the LHY correction to the GP equation, on the formation and dynamics of individual vortices and vortex lattices. The main outcome of our analysis is related to the scenario of vortex production when the only source of nonlinearity is provided by the LHY term. This condition may approximately hold for the single-component dipolar BEC, with the MF (mean-field) contact interaction and DDI (dipole-dipole interaction) roughly canceling each other. To achieve this setting, one has to select appropriate values of the strengths of the MF contact and dipole-dipole interactions, which can be provided through the Feshbach resonance, and by adjusting the angle that determines the orientation of the atomic magnetic moments. It is shown that the LHY-only nonlinearity can be made strong enough to produce vortices in the dipolar superfluid. For the smaller numbers of stable vortices ($N_{v}\leq 5$), it being more likely to obtain even rather than odd numbers of vortices when varying the rotational frequency $\Omega$. This means that the range of $\Omega$ is relatively large to generate even numbers ($2$ or $4$) of the vortices, being much smaller for the odd numbers ($1$ and $3$). This feature is particularly pronounced in the pure LHY case, so that, for the single-vortex case, the corresponding range of the rotation frequency $\Omega$ is so narrow that it can be identified with the critical one $\Omega_{c}$ for the production of the single vortex. This effect, related to the quartic LHY nonlinearity, can also be partially attributed to the aspect ratio of the 3D trap. In a less pronounced form, a similar effect was also noted for the relatively small values of vortices in the case of pure MF contact interactions. A more detailed investigation related to this effect, considering different parametrization, may be required to clarify the role of the aspect ratio and its interplay with the nonlinearity.
We have also computed, within the framework of our quasi-2D model, the number of vortices, chemical potential, and energy per particle, expressed as functions of the rotation frequency $\Omega$. For that purpose, we have considered specific combinations of the strengths of the MF contact interaction, DDI, and LHY nonlinearity. These parameters are assumed to be large enough to study the vortex production from the perspective of possible experimental realizations in dipolar BEC systems. The results demonstrate that, while the vortex numbers consistently increase with the growth of $\Omega$, the chemical potential and energy decrease. From this study, when considering the pure LHY nonlinearity, exemplified by a particular value of the strength, which can be further generalized, we conclude how to precisely identify the single-vortex critical frequency, along with the corresponding values of the chemical potential and energy. In all the investigated cases, we have observed results consistent with the Feynman’s prediction of linear dependence between angular momentum per particle and the number of vortices.
Acknowledgements.
S.S. thanks Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP) for the fellowship support, Proj. 2024/04174-8. R.R. wishes to acknowledge the financial assistance from DST-CURIE(DST-CURIE/PG/54) and ANRF( CRG/2023/008153) L.T. also acknowledges partial support from FAPESP (Proj. 2024/01533-7) and Conselho Nacional de Desenvolvimento Científico e Tecnológico (Proc. 303263/2025-3). The work of B.A.M. was supported, in part, by the Israel Science Foundation, through grant No. 1695/22.